the maximum current that can be measured by a galvanometer of resistance 40 ohm is 10mA. It is converted into a voltmeter that can read upto 50V. thg\e resistance to be connected in series with the galvanometer is

Grade:12th pass

2 Answers

Arun
5 years ago
To convert a galvanometer into voltmeter, the necessary value of resistance to be connected in series with the galvanometer is
R = V/ Ig - G
R = 5000 - 40
R = 4960 ohm
 
Regards
Arun (askIITians forum expert)
Gitanjali Rout
5 years ago
q]
To convert a galvanometer into voltmeter, the necessary value of resistance to be connected in series with the galvanometer is
R = V/ Ig  -  G
R = 5000 - 40
R = 4960 ohm